OpenStudy (anonymous):

Create your own piecewise function and graph it on the coordinate plane. Explain whether your piecewise function is continuous or discontinuous.

OpenStudy (anonymous):

\[f(x) = |x+4| = \left\{\begin{array}{rcc} x + 4 & \text{if} & x \geq -4 \\ - x - 4& \text{if} & x < -4 \end{array} \right. \] that one is continuous. but you can make any you like using this template from zarkon

OpenStudy (anonymous):

thank you :)